OpenStudy (anonymous):

Find the exact area of the circle with the given radius. r = 11.64 m A. 11.64π m² B. 23.28π m² C. 33.8724π m² D. 135.4896π m²

OpenStudy (anonymous):

a=π•r^2 this is the formula, I'm not sure why it doesn't work.

OpenStudy (anonymous):

I got 425.653132

OpenStudy (anonymous):

try 11.64 x 11.64

OpenStudy (anonymous):

oh, never mind haha, it's 135.4896πm^2

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

Find the exact area of the circle with the given diameter. d = 12.6 cm A. 39.69 cm2 B. 158.76 cm2 C. 158.76π cm2 D. 39.69π cm2

OpenStudy (anonymous):

i got c 158.76

OpenStudy (anonymous):

d/2=6.3 6.3^2•π=39.69πcm^2

OpenStudy (anonymous):

Which is the closest approximation to the area of the circle? Use 3.14 or to approximate π. A. 531 ft² B. 169 ft² C. 82 ft² D. 41 ft²

OpenStudy (anonymous):

13^2•π=530.929≈531

OpenStudy (anonymous):

What is the exact area of a circular courtyard with a radius of 22 ft? A. 1936π ft² B. 484π ft² C. 44π ft² D. 22π ft²

OpenStudy (anonymous):

b?

OpenStudy (anonymous):

yes it's B:)

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

Get Lucky - Daft Punk (OFFICIAL RADIO EDIT)

OpenStudy (anonymous):

nice song

OpenStudy (anonymous):

oh have you heard of havent met you yet?

OpenStudy (anonymous):

Which expression can be used to find the area of a circle with a diameter of 80 cm? A. π · 402 B. π · 160 C. π · 80 D. π · 40

OpenStudy (anonymous):

thats my favorite songs:)

OpenStudy (anonymous):

really

OpenStudy (anonymous):

Michael Bublé - "Haven't Met You Yet" [Official Music Video]

OpenStudy (anonymous):

thats my favorite song it's funny when they were walking on the racks

OpenStudy (anonymous):

80/2=40=r 40^2•π=a is A π•40^2?

OpenStudy (anonymous):

im a fan of michael buble

OpenStudy (anonymous):

same here you have a skype?

OpenStudy (anonymous):

Omgggg i got a 100%%%%%%%

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

yay!!!

OpenStudy (anonymous):

ima add you on skype

OpenStudy (anonymous):

haha ok